paige12345:

In order to solve for the variable in the equation 2 (x + 3) + 5 x = 3 (2 x minus 1), Jaleesa begins by applying the distributive property, then combines like terms. Which equation is the result of these steps?

BestBoiKirby:

is this a test q?

mw104969:

7x+6=6x+3

BestBoiKirby:

is this a test q tho?

paige12345:

q

mw104969:

is that what you were looking for^^

paige12345:

no its not one of the choses.

mw104969:

Show the choices

paige12345:

In order to solve for the variable in the equation 2 (x + 3) + 5 x = 3 (2 x minus 1), Jaleesa begins by applying the distributive property, then combines like terms. Which equation is the result of these steps? 7 x + 6 = 6 x minus 3 2 x + 6 + 5 x = 6 x minus 3 2 x + 3 + 5 x = 6 x minus 1 7 x + 3 = 6 x minus 1

mw104969:

its the first choice